OpenStudy (anonymous):

solve the system of equations and state the answer as an ordered pair. x+y=-5 x+y=6 3x+4y=2 2x+3y=1

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@james08

OpenStudy (rhystic):

the first one is not a proper system of equations.. by definition if x+y = -5 and x+y = 6 then -5 should equal 6 which it does not. It is not solveable. The second you can just just use substition. solve for y in the first which yields: 3x+4y = 2 -> y = -(3/2)x + 1/2 and sub it in for y in the second equation and solve for x then go back and solve for y.